OpenStudy (anonymous):

Which of the following correctly uses absolute value to show the distance between –70 and 13? |–70 – 13| = |–83| = –83 units |–70 – 13| = |–83| = 83 units |–70 + 13| = |–57| = 57 units |–70 + 13| = |–57| = –57 units

OpenStudy (anonymous):

@Mokeira

OpenStudy (anonymous):

|–70 – 13| = |–83| = 83 units

OpenStudy (anonymous):

|–70 + 13| = |–57| = 57 units

OpenStudy (anonymous):

what ? @yoonjae

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

@aryandecoolest

OpenStudy (mokeira):

@Dobby1

OpenStudy (dobby1):

yeah?

OpenStudy (anonymous):

help lol

OpenStudy (mokeira):

do you know how to go about the question above?

OpenStudy (dobby1):

let me take a look I might be able to simplify it more

OpenStudy (anonymous):

thanks ^

OpenStudy (dobby1):

It is |–70 + 13| = |–57| = 57 units This is because this is the only one of those equations where the ending result is the actual distance between -70 and 13 which is 57 units

OpenStudy (dobby1):

I know I gave the answer but it was really the only way to help you here so I made sure to explain it afterwards but I wish i could helped you find it your self

OpenStudy (kropot72):

To find the distance between the points, the values at the points must be subtracted: Distance = |-70 - 13| = |-83| = 83 units

OpenStudy (vickyisthesmartone123):

go to mathway.com type in the equation choose what you want and BUALA answer is their bye!